# Conformal Validity Guarantees Exist for Any Data Distribution

### v2

@@ B.2 Deriving Practical Conformal Algorithms under Agent-Induced MFCS
Appendix C Details on the Roles of Exchangeability, Weighted Exchangeability, and Pseudo-Exchangeability in Conformal Prediction
C.1 The Role of Exchangeability in Conformal Prediction
@@
Appendix A.6 of Tibshirani et al. (2019) provides an alternate proof to a key Lemma underlying (standard) conformal prediction’s validity, which essentially presents conformal prediction as an inverted permutation test (though they do not explicitly state this). We reproduce most of that argument here, with some additional explanatory commentary to highlight the role played by the exchangeability assumption in standard conformal prediction.
The general strategy in Tibshirani et al. (2019) is “to condition on the unlabeled multiset of values $\{v_{1},...,v_{n+1}\}$ obtained by our random variables $V_{1},...,V_{n+1}$, and then inspect the probabilities that the last random variable $V_{n+1}$ attains each one of these values.” That is, Tibshirani et al. (2019) condition on the event $\{V_{1},...,V_{n+1}\}=\{v_{1},...,v_{n+1}\}$, which they denote by $E_{v}$—importantly, this is not saying that $V_{i}=v_{i}$ for all $i\in\{1,...,n+1\}$. Rather, it is only saying that the set or “bag” [^14] For simplicity, Tibshirani et al. (2019) assume there are almost surely no ties among those scores (to work with sets rather than multisets/bags); the argument applies in the general case but with more cumbersome notation. of values $\{v_{1},...,v_{n+1}\}$ are the values obtained by the random variables $V_{1},...,V_{n+1}$; we do not yet know whether $v_{1}$ is the value obtained by $V_{1}$, or by $V_{2}$, or by $V_{n+1}$, etc.
@@
This statement can be thought of as an inverted permutation test as it equates the probability $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}$ to the total probability mass of all permutations $\sigma$ that map $n+1$ to $i$, normalized across all possible permutations of the $n+1$ values. A more explicit way of writing this could include the random variables in the joint density function’s notation, to emphasize that the permutation is on the values and not on the random variables:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}=\frac{\sum_{\sigma:\sigma(n+1)=i}f(V_{1}=v_{\sigma(1)},...,V_{n+1}=v_{\sigma(n+1)})}{\sum_{\sigma}f(V_{1}=v_{\sigma(1)},...,V_{n+1}=v_{\sigma(n+1)})}.$
Note that each argument in $f$ still corresponds to the index of the random variable (not the value); meanwhile, the positions of the observed values are permuted to argument locations.
Next, the key idea is then that—due to exchangeability—every permutation of the data is equally likely, and so the probability of the event $\{V_{n+1}=v_{i}\ |\ E_{v}\}$ can be computed as the total number of permutations (of the $n+1$ values) where $\sigma(n+1)=i$, divided by the total number of all possible permutations (of the $n+1$ values). I.e., by exchangeability $f(V_{1}=v_{\sigma(1)},...,V_{n+1}=v_{\sigma(n+1)})=f(V_{1}=v_{1},...,V_{n+1}=v_{n+1})$ for all $\sigma$, so we can simplify to reduce our statement to counting permutations:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}$ $=\frac{\sum_{\sigma:\sigma(n+1)=i}f(V_{1}=v_{1},...,V_{n+1}=v_{n+1})}{\sum_{\sigma}f(V_{1}=v_{1},...,V_{n+1}=v_{n+1})}$
@@
That is, the third line follows from the second because (in the denominator’s summation) there are $(n+1)!$ possible permutations of $n+1$ values and (in the numerator’s summation) there are $n!$ possible permutations of $n$ values (since in the numerator’s summation we only consider $\sigma$ where $\sigma(n+1)=1$, there are only $n$ values unknown).
Tibshirani et al. (2019) then explain that this result shows that the distribution of $V_{n+1}|E_{z}$ is uniform on the set $\{v_{1},...,v_{n+1}\}$, that is,
@@
where $\delta_{v}$ denotes a point mass at the value $v$. The remaining steps of the proof then follow by the definition of a quantile, by the meaning of conditioning on $E_{z}$, and by marginalizing.
The role of exchangeability is to allow us to count every permutation equally in the inverted permutation test. In the next sections, we will see that under weighted exchangeability, we instead need to use the weight functions to see how much “importance” to place on each of the permutations.
C.2 The Role of Weighted Exchangeability in Conformal Prediction

### v3

@@ B.2 Deriving Practical Conformal Algorithms under Agent-Induced MFCS
Code for our specific implementations of MFCS Full CP and Split CP methods are available at the following GitHub repository: https://github.com/drewprinster/conformal-mfcs.
B.3 Alternative Approaches to Deriving MFCS Conformal Guarantees and Algorithms
Here we describe certain alternative approaches to the derivation steps in Appendix B.1, and we briefly discuss how these alternate steps could have implications for practical algorithm design for MFCS settings and more broadly.
Alternative Step 1: Assumptions via a Probabilistic Graphical Model
Rather than beginning with the MFCS conditional independence assumptions (Eq. (3)) for Step 1, we could have instead begun by describing our data-generating process using a probabilistic graphical model (Koller & Friedman, 2009; Pearl, 2009) and used the conditional independence conditions implied by the missing edges in that model. [^17] Missing edges in probabilistic graphical models imply conditional independences in the joint PDF $f$; the converse depends on further assumptions (Pearl, 2009). One advantage of graphical models is that they can be more visually intuitive than conditional independence or invariance assumptions. For example, consider the directed acyclic graph (DAG) model in Figure 4, on which we take a causal interpretation to view each directed edge $A\rightarrow B$ as representing that $A$ has a causal effect on $B$ [^18] We take posit a causal DAG here so that the edges better map onto our intuition about cause and effect; however, in other settings different graphical models could instead be used so long as the appropriate Markov property is used to obtain the implied conditional independence relations. As in MFCS, in this causal DAG model we first assume $Z_{1:n}=\{Z_{1},...,Z_{n}\}$ are independent and identically distributed (IID), and in the DAG we represent these IID variables with a single node to be concise; for the remaining nodes, we allow each $X_{n+t}$ to depend causally on all prior observations (i.e., each $X_{n+t}$ has incoming edges from all prior nodes), while we only allow each $Y_{n+t}$ to be affected by $X_{n+t}$ (i.e., $Y_{n+t}$ only has an incoming edge from $X_{n+t}$). We use blue arrows to represent relationships that we further assume are equivalent (in distribution).
[…]
Figure 4: A causal directed acyclic graph (DAG) model that implies $Z_{1},...,Z_{n+t-1}\perp\!\!\!\perp Y_{n+t}\mid X_{n+t}$ (as does MFCS). Moreover, the blue edges represent relationships that are further assumed to be equivalent, implying $Y_{n+t}|X_{n+t}\stackrel{{\scriptstyle\mathclap{d}}}{{=}}Y|X$ for all $t\in\{1,...,T\}$.
By the rules of $d$-separation for DAGs (Pearl, 1988), for any $t\in\{1,...,T\}$, conditioning on $X_{n+t}$ blocks all paths from prior nodes to $Y_{n+t}$; so, this DAG implies the same conditional independence relation that we obtained from our original MFCS invariance assumptions: $Z_{1},...,Z_{n+t-1}\perp\!\!\!\perp Y_{n+t}\mid X_{n+t}$. Moreover, the blue edges represent relationships that are further assumed to be equivalent in distribution, implying $Y_{n+t}|X_{n+t}\stackrel{{\scriptstyle\mathclap{d}}}{{=}}Y|X$ for all $t\in\{1,...,T\}$.
Alternative Steps 2 and 3: Factoring $f$ with likelihood-ratio factors
The following is an alternative approach to Steps 2 and 3 in Appendix B.1, which factorizes $f$ using likelihood-ratio factors, which more closely resemble the specific CP algorithms for standard covariate shift in Tibshirani et al. (2019) and (one-step) feedback covariate shift in Fannjiang et al. (2022). We will refer to the factorization given in Step 2 of Appendix B.1 as the direct likelihood factorization, and the alternative derivation we present here as the likelihood ratio-based factorization; we will soon discuss tradeoffs of these different viewpoints regarding interpretability and implementation.
First, recall the result of Step 2 in Appendix B.1:
$f$ $(z_{1},...,z_{n+t})=\prod_{j=1}^{n+t}\Big{[}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\mid z_{1},...,z_{j-1})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{j}|x_{j})\Big{]}.$
Whereas the direct factorization in Appendix B.1 proceeds immediately from here to Step 3 by plugging this statement into Eq. (4), now we consider further steps to obtain explicit likelihood-ratio factors. To do so, we simplify using our assumption that the first $n$ points are IID, which implies $p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\ |\ z_{1},...,z_{j-1})=p_{X}(x_{j})$ for $j\in\{1,...,n\}$, and then use a “multiply by one” trick to multiply by $1=\tfrac{\prod_{j=n+1}^{n+t}p_{X}(x_{j})}{\prod_{j=n+1}^{n+t}p_{X}(x_{j})}$, where $p_{X}:=dP_{X}^{(0)}$ denotes the density for the features in the initial IID distribution $P_{Z}^{(0)}=P_{X}^{(0)}\times P_{Y|X}$:
$f(z_{1},...,z_{n+t})$ $=\prod_{j=1}^{n}\Big{[}p_{X}(x_{j})\Big{]}\cdot\prod_{j=n+1}^{n+t}\Big{[}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\ |\ z_{1},...,z_{j-1})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{j}|x_{j})\Big{]}$
$=\prod_{j=1}^{n}\Big{[}p_{X}(x_{j})\Big{]}\cdot\prod_{j=n+1}^{n+t}\Big{[}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\ |\ z_{1},...,z_{j-1})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{j}|x_{j})\Big{]}\cdot\frac{\prod_{j=n+1}^{n+t}p_{X}(x_{j})}{\prod_{j=n+1}^{n+t}p_{X}(x_{j})}$
$=\prod_{j=1}^{n+t}\Big{[}p_{X}(x_{j})\Big{]}\cdot\prod_{j=n+1}^{n+t}\Big{[}\frac{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\ |\ z_{1},...,z_{j-1})}{p_{X}(x_{j})}\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{j}|x_{j})\Big{]}$
$=\prod_{j=n+1}^{n+t}\Big{[}\frac{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\ |\ z_{1},...,z_{j-1})}{p_{X}(x_{j})}\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{j}|x_{j})\cdot p_{X}(x_{j})\Big{]}$
$=\prod_{j=n+1}^{n+t}\Big{[}\underbrace{\frac{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\ |\ z_{1},...,z_{j-1})}{p_{X}(x_{j})}}_{\begin{matrix}\text{Time-dependent }\\ \text{likelihood-ratio factors}\end{matrix}}\Big{]}\cdot\underbrace{\prod_{j=1}^{n+t}\Big{[}p_{Z}(z_{j})\Big{]}}_{\begin{matrix}\text{Time-invariant}\\ \text{``core'' function }\end{matrix}}.$
There are two main tradeoffs between the direct likelihood factorization (Step 2 of Appendix B.1) and the alternative likelihood-ratio factorization presented here—the first tradeoff is with respect to interpretability, and the second tradeoff is with respect to implementation. First, regarding interpretability: a benefit of the factorization presented in Appendix B.1 is that its permutation-invariant factor $\prod_{j=1}^{n+t}p_{Y|X}(y_{j}\mid x_{j})$ clearly and intuitively corresponds to the assumption that $Y\mid X$ is invariant in MFCS, which is one reason we chose to use that factorization in our main paper’s presentation; on the other hand, a benefit of the current likelihood-ratio factorization is that it can be interpreted more clearly as “the likelihood obtained by adjusting relative to if all the data were IID,” since the permutation-invariant factor $\prod_{j=1}^{n+t}p_{Z}(z_{j})$ represents the probability of the sequence of observations if they were all IID, and the likelihood ratios represent the necessary adjustment factors needed to compute the true probability. The current likelihood ratio-based factorization also relates more closely to the analysis and algorithms in Tibshirani et al. (2019) and Fannjiang et al. (2022).
Next, to discuss implementation tradeoffs, it may be helpful to first complete Step 3 by plugging our factorization into the equation (Eq. (4)) for the CP weights:
$\frac{\sum_{\sigma:\sigma(n+t)=i}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}$ $=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=n+1}^{n+t}\Big{[}\tfrac{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}{p_{X}(x_{\sigma(j)})}\Big{]}\prod_{j=1}^{n+1}\Big{[}p_{Z}(z_{\sigma(j)})\Big{]}}{\sum_{\sigma}\prod_{j=n+1}^{n+t}\Big{[}\tfrac{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}{p_{X}(x_{\sigma(j)})}\Big{]}\prod_{j=1}^{n+1}\Big{[}p_{Z}(z_{\sigma(j)})\Big{]}}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=n+1}^{n+t}\Big{[}\tfrac{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}{p_{X}(x_{\sigma(j)})}\Big{]}}{\sum_{\sigma}\prod_{j=n+1}^{n+t}\Big{[}\tfrac{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}{p_{X}(x_{\sigma(j)})}\Big{]}}.$
Completing this Step 3 emphasizes that it is the “dynamic” factors (that are not permutation invariant) which are of practical concern for implementation, since the permutation-invariant factor $\prod_{j=1}^{n+t}p_{Z}(z_{\sigma(j)})$ has cancelled out in the ratio.
One implementation tradeoff between the direct (Appendix B.1) and the likelihood ratio-based factorization here is an epistemic one: Implementations using the direct factorization rely on computing or estimating the product of likelihoods $\prod_{j=1}^{n+t}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})$; we leverage this direct approach in our main paper largely because our practical experiments focus on ML-agent-induced MFCS where $p_{X_{j}|Z_{1},...,Z_{j-1}}(x\mid Z_{1},...,Z_{j-1})$ is known (it is the ML agent’s query function). On the other hand, for a hypothetical MFCS instance where $p_{X_{j}|Z_{1},...,Z_{j-1}}(x\mid Z_{1},...,Z_{j-1})$ is not known, an implementation based on the likelihood-ratio factorization may be preferred, because likelihood-ratio (density-ratio) estimation can often be easier than direct likelihood (density) estimation.
Another implementation consideration to note is the “failure mode” of specific permutations $\sigma$ that have zero probability density, for example due to $p_{X}(x_{\sigma(j)})=0$ for some $j\in\{n+1,...,n+t\}$, which could occur if $X_{n+t}=x_{n+t}$ is observed out of the support of $P_{X}^{(0)}$, the initial distribution that the first $n$ points are assumed to be drawn IID from in MFCS. Implementations based on the likelihood ratio-based factorization would be undefined in such cases due to requiring dividing by zero, and thus they may need to exclude these cases by assuming that the density in the numerator is absolutely continuous with respect to that of the denominator (e.g., Tibshirani et al. (2019) and Fannjiang et al. (2022) made this assumption for their specific CP algorithms for one-step standard and feedback covariate shift). On the other hand, implementations using the direct factorization (Appendix B.1) are not necessarily undefined in these cases—that is, if one commits to a definition of conditional probability densities where $p(A|B)=0$ if $p(B)=0$, rather allowing the conditional density to be undefined in such cases (see e.g., Kolmogorov (1956)). For a specific example, let us assume $T=1$ and that $X_{n+1}=x_{n+1}$ is observed out of the support of $P_{X}^{(0)}$, so $p_{X}(x_{n+1})=0$. Then (assuming a conditional probability density definition where $p(A|B)=0$ if $p(B)=0$), the MFCS weights obtained via direct factorization in Eq. (14) would reduce to $\mathbb{P}_{n+1}^{(\text{MFCS})}\{z_{i}\mid E_{z}^{(t)}\}=0$ for all $i\neq n+1$, and $\mathbb{P}_{n+1}^{(\text{MFCS})}\{z_{n+1}\mid E_{z}^{(t)}\}=1$. Accordingly, the resulting CP set would maintain coverage trivially by concentrating all the weight on $\delta_{\infty}$ (the point mass at $\infty$) in the (conservative) empirical score distribution, thus forcing a non-informative prediction set $\widehat{C}_{n}(x)=\mathcal{Y}$.
Appendix C Bounded Query Function Experiments for Adaptive Exploration
C.1 Bounding the Query Function to Enforce Informative (Finite) Prediction Intervals
Because noninformative prediction sets ($\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$) occur in our weighted CP MFCS experiments when a test point’s normalized weight as in Eq. (9) (implemented as in Eq. (16)) exceeds $\alpha$, we now consider redesigning the ML query probability functions to prevent this from happening. For simplicity, we impose this constraint on the one-step weights. In particular, we consider bounding the query probability function $p(x\mid Z_{1},...,Z_{j-1})$ so that instead of being proportional to $\exp(\lambda\cdot u_{t}(x))$, the probabilities are instead proportional to a bounded function:
$p(x\mid Z_{1},...,Z_{j-1})\propto$ $\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)},\quad\text{where }B=\sup\bigg{\{}b>0:\frac{b}{\sum_{i=1}^{n}\min\big{(}\exp(\lambda\cdot u_{t}(X_{i})),b\big{)}+b}<\alpha\bigg{\}}.$
That is, we select $B$ as the largest bound that we could use to avoid infinite-width prediction intervals (in the depth $d=1$ weights). The selection of $B$ above avoids noninformative prediction sets because
$\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)}\leq B$
In our implementations, we solve for $B$ via a binary search algorithm over all values of the unbounded function $\exp(\lambda\cdot u_{t}(x))$ for data in the input space $x\in\mathcal{X}$, which is tractable in our experiments where $\mathcal{X}$ is the finite pool or library of possible query points. That is, we specifically solve for $B=\max\big{\{}b=\exp(\lambda\cdot u_{t}(x))\ :\ x\in\mathcal{X},\ \tfrac{b}{\sum_{i=1}^{n}\min(\exp(\lambda\cdot u_{t}(X_{i})),b)+b}<\alpha\big{\}}$. Code for our implementations is available here: https://github.com/drewprinster/conformal-mfcs.
Across the four real datasets and the same experimental conditions we used for our active learning experiments in Figure 3 of our main paper, the following results empirically validate that imposing the proposed adaptive bound on our MFCS CP method’s query function prevents any noninformative (infinite width) intervals, all without sacrificing coverage even over a long time horizon (70 active learning steps). For easy comparison, for each dataset we provide the corresponding results for the original unbounded query function from Figure 3. We moreover plot the “relative magnitude” of the bound $B$—i.e., we plot $B/B_{\max}$, where $B_{\max}=\max_{x\in\mathcal{X}}(\exp(\lambda\cdot\hat{\sigma}(x)))$—across the same trajectory of active learning steps.
Airfoil dataset
Figure 5: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the airfoil dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds. Hyperparameters for the experiments are given in Appendix E.2 Table 3.
Communities dataset
Figure 6: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the communities dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds. Hyperparameters for the experiments are given in Appendix E.2 Table 3.
Medical Expenditure Panel Survey (MEPS) dataset
Figure 7: Unbounded (top row) versus bounded (bottom row) active learning experiments with proposed multistep split CP method for estimation depth $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the Medical Expenditure Panel Survey (MEPS) dataset. The Y-axes represent mean coverage, median interval width, mean squared error on a holdout test set, and the relative bound magnitude $B/B_{\text{max}}$; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds. Hyperparameters for the experiments are given in Appendix E.2 Table 3.
Blog dataset
Figure 8: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the blog dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds. Hyperparameters for the experiments are given in Appendix E.2 Table 3.
The bounded query function’s improvement in prediction-interval informativeness (smaller prediction intervals relative to the unbounded query function) sometimes comes at the cost of an initially delayed decrease in MSE, though the long-run MSE for the bounded and unbounded query functions are comparable for most datasets. The bounded query function can thus be thought of as an “informativeness constraint” that forces the AI/ML agent to initially be more “cautious” or “slow” in its exploration until it has accumulated enough training and calibration data to explore more aggressively (as evidenced by the increasing bound magnitude over time).
@@ C.1 Bounding the Query Function to Enforce Informative (Finite) Prediction Intervals
Appendix D Details on the Roles of Exchangeability, Weighted Exchangeability, and Pseudo-Exchangeability in Conformal Prediction
D.1 The Role of Exchangeability in Conformal Prediction
@@
Appendix A.6 of Tibshirani et al. (2019) provides an alternate proof to a key Lemma underlying (standard) conformal prediction’s validity, which corresponds to viewing conformal prediction as an inverted permutation test (though they do not explicitly state this). We reproduce most of that argument here, with some additional explanatory commentary to highlight the role played by the exchangeability assumption in standard conformal prediction.
The general strategy in Tibshirani et al. (2019) is “to condition on the unlabeled multiset of values $\{v_{1},...,v_{n+1}\}$ obtained by our random variables $V_{1},...,V_{n+1}$, and then inspect the probabilities that the last random variable $V_{n+1}$ attains each one of these values.” That is, Tibshirani et al. (2019) condition on the event $\{V_{1},...,V_{n+1}\}=\{v_{1},...,v_{n+1}\}$, which they denote by $E_{v}$—recall that this is not saying that $V_{i}=v_{i}$ for all $i\in\{1,...,n+1\}$. Rather, it is only saying that the set or “bag” [^19] For simplicity, Tibshirani et al. (2019) assume there are almost surely no ties among those scores (to work with sets rather than multisets/bags); the argument applies in the general case but with more cumbersome notation. of values $\{v_{1},...,v_{n+1}\}$ are the values obtained by the random variables $V_{1},...,V_{n+1}$; we do not yet know whether $v_{1}$ is the value obtained by $V_{1}$, or by $V_{2}$, or by $V_{n+1}$, etc.
@@
This statement can be thought of as an inverted permutation test as it equates the probability $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}$ to the total probability mass of all permutations $\sigma$ that map $n+1$ to $i$, normalized across all possible permutations of the $n+1$ values. For intuition, more explicit notation with random variables can be helpful: That is, if $f$ is a PDF, then writing $f(v_{\sigma(1)},...,v_{\sigma(n+1)})=f_{V_{1},...,V_{n+1}}(v_{\sigma(1)},...,v_{\sigma(n+1)})$; if $f$ is a PMF, then writing $f(v_{\sigma(1)},...,v_{\sigma(n+1)})=f(V_{1}=v_{\sigma(1)},...,V_{n+1}=v_{\sigma(n+1)})$. In either case, writing the random variables explicitly can help to emphasize that each argument in $f$ still corresponds to the index of the random variable; meanwhile, the observed values are being permuted to different argument locations.
Next is where exchangeability comes in. That is, exchangeability means $f(v_{\sigma(1)},...,v_{\sigma(n+1)})=f(v_{1},...,v_{n+1})$ for permutations $\sigma$ of the indices $\{1,...,n+1\}$, so by plugging in this substitution and factoring out $f(v_{1},...,v_{n+1})$ from the summation, we can simplify to reduce our statement to counting permutations:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}$ $=\frac{\sum_{\sigma:\sigma(n+1)=i}f(v_{1},...,v_{n+1})}{\sum_{\sigma}f(v_{1},...,v_{n+1})}$
$=\frac{f(v_{1},...,v_{n+1})\sum_{\sigma:\sigma(n+1)=i}1}{f(v_{1},...,v_{n+1})\sum_{\sigma}1}$
@@
That is, the last line follows because (in the denominator’s summation) there are $(n+1)!$ possible permutations of $n+1$ values and (in the numerator’s summation) there are $n!$ possible permutations of $n$ values (since in the numerator’s summation we only consider $\sigma$ where $\sigma(n+1)=1$, there are only $n$ values unknown). In other words, under the assumption of exchangeability, every permutation of the data is equally likely, so the probability of the event $\{V_{n+1}=v_{i}\ |\ E_{v}\}$ can be computed as the total number of permutations (of the $n+1$ values) where $\sigma(n+1)=i$, divided by the total number of all possible permutations (of the $n+1$ values).
The remainder of the argument in Tibshirani et al. (2019) then follows similarly as in the proof given in Appendix A.1, but premised on the assumption of exchangeability. That is, Tibshirani et al. (2019) explain that the above result shows that the distribution of $V_{n+1}|E_{z}$ is uniform on the set $\{v_{1},...,v_{n+1}\}$:
@@
where $\delta_{v}$ denotes a point mass at the value $v$. The remaining steps of the proof then follow by the definition of a quantile, by the meaning of conditioning on $E_{z}$, and by marginalizing (similarly as in Appendix A.1). The resulting standard CP coverage guarantee can be viewed as a corollary to Theorem 4.1 because its proof (as presented by Tibshirani et al. (2019)) follows the same steps, except also relying on the simplifying assumption of exchangeability to allow us to count every permutation equally.
Lastly, we note that while standard conformal prediction is often defined by taking the level $(1-\alpha)\cdot\tfrac{n+1}{n}$ quantile over the values $\{v_{1},...,v_{n}\}$, this is equivalent to taking the level $1-\alpha$ quantile of values $\{v_{1},...,v_{n}\}\cup\{\infty\}$. Tibshirani et al. (2019) introduced the latter with the point mass $\delta$ notation we have been using to enable writing a weighted empirical distribution of score values.
D.2 The Role of Weighted Exchangeability in Conformal Prediction